NvidiaIntely (nvidiaintely):

To get from your house to your grandparent’s house you must avoid walking through a pond. To avoid the pond, you walk 0.5 miles south and 1.2 miles east. How many miles would be saved if it were possible to walk through the pond? 0.4 miles 1.69 miles 3.0 miles 6.4 miles

OpenStudy (kevin):

x^2 = 1.2^2 + 0.5^2 x = 1.3 saved --> 1.2+0.5-1.3 = 0.4

NvidiaIntely (nvidiaintely):

Thanks!